# A Theory of Formalisms for Representing Knowledge

### v1

@@ Introduction
Knowledge plays a crucial role in a large number of fields such as intelligent decision making, natural language understanding and automatic scientific discovery. How to effectively represent, acquire, utilize and evolve knowledge is undoubtedly one of the most critical parts of realizing artificial general intelligence (AGI). Knowledge representation is the starting point and foundation of all of these tasks. In traditional knowledge representation and reasoning (KR), representations of knowledge are often regarded as “explicit, symbolic, declarative representations of information” (Delgrande et al. 2024). However, in this work, we aim to explore more general forms of knowledge representation.
Over the past seven decades, researchers have devoted substantial efforts to developing various knowledge representation formalisms. A far incomplete list includes: monotonic logical systems such as Prolog (van Emden and Kowalski 1976) and description logics (Baader et al. 2017); non-monotonic logics such as circumscription (McCarthy 1980) and default logic (Reiter 1980); graph-based representations such as Bayesian networks (Pearl 1985) and semantic network (Sowa 1991), and parametrized models such as recurrent neural networks (Rumelhart, Hinton, and Williams 1986) and transformers (Vaswani et al. 2017). It is noteworthy that machine learning and knowledge representation are inherently intertwined; all learning algorithms are actually based on some formalisms for representing knowledge.
The quest for the best formalism of knowledge representation has sparked a longstanding dispute. A prime example is the “declarative vs. procedural” controversy, which centers on choosing between declarative statements and procedures as the primary means of representing knowledge. Similarly, the ongoing debate between symbolic AI and connectionist AI revolves around the question of whether knowledge should be represented explicitly or implicitly. In general, knowledge representation formalisms in machine learning, such as convolutional neural networks in deep learning and transformers in large language models, are usually implicit, while all logical formalisms in traditional KR are explicit.
@@
The main contributions of this work are threefold. Firstly, we propose a general framework to capture all the knowledge representation formalisms in which we are interested, and propose a novel definitions for universal formalisms. Secondly, we find a family of universal formalisms; based on them, we then prove that all possible universal formalisms are recursively isomorphic. Thirdly, we show that, under a natural condition, all subrecursive formalisms that can be translated into each other are recursively isomorphic. These results show us that, if an offline compilation is allowed, almost all the natural representation formalisms with the same expressive power can be regarded as the same, which thus provides us a partial answer to the aforementioned dispute. [^1] Due to space limitations, many proofs have been omitted from the main text. Additional details can be found in the appendix.
@@
Given a subclass $K$ of $\mathscr{D}\times\mathscr{Q}$, let $\textit{cl}(K)$ denote the minimum superclass of $K$ which admits Properties 1-5 in Definition 3. We need to assure the desired transformation $(\cdot)^{\ast}$ satisfying the property: $\mathbb{K}(M^{\ast})=\textit{cl}(\mathbb{K}(M))$.
@@
To address these challenges, we introduce a task array $T$, where the $i$-th task is stored in $T[i]$. Although there could be an infinite number of nonterminating tasks in $T$, we can apply a standard technique to sequentially simulate multiple such tasks. This can be visualized using a coordinate system where the x-axis represents the number of steps and the y-axis denotes task indices. The simulation is actually a procedure to traverse the first quadrant of this coordinate system.
An atomic task is a tuple $t:=(N,p_{1},\dots,p_{k})$, where $N$ is a Turing machine and $p_{1},\dots,p_{k}$ its parameters. Performing $t$ involves simulating $N$ on the input $[\![p_{1},\dots,p_{k}]\!]$, with $t$ being successful if $N$ accepts. Every task is either an atomic task or a finite sequence of atomic tasks $t:=\langle t_{1},\dots,t_{n}\rangle$. Executing $t$ means sequentially performing the atomic tasks $t_{1},\dots,t_{n}$. The task $t$ is said to be successful or to succeed if all its constituent atomic tasks succeed.
Moreover, let $M_{e}$ denote a Turing machine that accepts $[\![\phi,\psi]\!]$ iff $\phi,\psi$ are a pair of FO-sentences such that $\phi\vDash\psi$. For a technical reason, let $M_{n}$ denote a Turing machine that accepts $[\![\tau]\!]$ iff $\tau$ is an injective partial functions on $\Delta$.
@@
to $T$, where $\chi$ is the next query after $\psi$ in $\mathscr{Q}$. By repeating this process, we effectively enumerate all queries in $\mathscr{Q}$. For details, see Lines 22-23, 27-28, and 42-47 in Procedure 1.
@@
A flag array $F$ records the completion status of each task as either true or false. The relationships between tasks are tracked using arrays $p$ (parent) and $c$ (cooperation). Specifically, $p[n]=i$ indicates that the task $T[n]$ is generated from the task $T[i]$, making $T[i]$ the parent of $T[n]$. A value of $c[i]=-1$ signifies that the task $T[i]$ operates independently. If $T[i]$ succeeds, its parent task’s flag, $F[p[i]]$, is set to true. Otherwise, if $c[i]=j\neq-1$, the task $T[i]$ must cooperate with the task $T[j]$. Only when both $T[i]$ and $T[j]$ succeed does the parent task’s flag, $F[p[i]]$, get set to true. For details on truth propagation, refer to Lines 14-16 in Procedure 1.
@@ 5 $n\leftarrow 1$; /* There is one task in the current array */
6 for $k\leftarrow 1$ to $\infty$ do
7       for $i\leftarrow 0$ to $\min(n,k)$ do
@@ 16                         $F[j]\leftarrow$ true;
17                  /* Accept if a support has been found */
18                   if $F[0]$ then accept;
19
@@ 58                   $n\leftarrow n+1$;
Let $\Theta$ denote the mapping which maps $[\![M^{\ast}]\!]$ to $\mathbb{K}(M^{\ast})$ for all Turing machines $M$.
@@
The following is a KRF-version of the recursion theorem.
@@
Now we are in the position to present the main theorem.
@@
Every FO-sentence $\phi$ is logical equivalent to $\phi\wedge\phi$. The defined KBs are thus the same, which provides a way to effectively find theories specified to the same KB.
@@ Appendix A Appendix: Detailed Proofs
Take $\gamma_{D_{0},\psi}\in\Sigma$ arbitrarily. If there exists no injective assignment $s$ in $\mathcal{B}$ such that $\mathcal{B}\models\theta_{D_{0}}^{v}[s]$, then $\mathcal{B}$ is trivially a model of $\gamma_{D_{0},\psi}$. Otherwise, let $s$ be any of such assignments. Let $\tau$ be a mapping that maps each $c\in\Delta$ to $s(v_{c})$. It is not difficult to verify that $\tau$ is injective and $D$ extends $\tau(D_{0})$. By the construction of $\Sigma$ and $\gamma_{D_{0},\psi}\in\Sigma$, we know $(D_{0},\psi)\in K$. Since $\mathscr{Q}$ is closed under constant renaming, it must be true that $\tau(\psi)\in\mathscr{Q}$. According to Property 5 in Definition 3, we thus have $(\tau(D_{0}),\tau(\psi))\in K$, and consequently, $\tau(\psi)\in\Psi$. This implies that $\mathcal{B}$ is a model of $\tau(\psi)$, or equivalently, $\mathcal{B}\models\psi[s]$. As a consequence, $\mathcal{B}$ is also a model of $\gamma_{D_{0},\psi}$ in this case. By the arbitrariness of $\gamma_{D_{0},\psi}$, we then conclude that $\mathcal{B}$ is a model of $\Sigma$. By definition, it is easy to see that $\mathcal{B}|_{\sigma_{Q}}=\mathcal{A}$. Consequently, we have $\mathcal{A}\models\phi$. This thus proves $\Psi\vDash_{u}\phi$ that we need.
@@
For convenience, we say a task in $T$ is render as true if the flag in $F$ corresponding to the the task is set to true.
We first consider the direction of $\subseteq$”. Let $(D,\phi)\in\mathbb{K}(M^{\ast})$. Now our task is to prove that $(D,\phi)\in\textit{cl}(\mathbb{K}(M))$. By definition, $M^{\ast}$ will accept the input $[\![D,\phi]\!]$. Let $N(D,\phi)$ denote the number of iterations of the while-loop (see Lines 14-16 in Procedure 1) before the task in which $(M,D,\phi)$ appears is rendered as true.
@@
We prove this by a routine induction on the length $n$ of $S$. For the case of $n=0$, $S$ consists of exactly one pair $(D,\phi)$, and we have that either $M$ accepts $[\![D,\phi]\!]$ or $\phi$ is a tautology. Thus, $t_{0}=(M,D,\phi)$ if the first case happens, and $t_{0}=(M_{e},\top,\phi)$ otherwise. It is easy to see that in either case performing $t_{0}$ can be done in a finite time. (Note that the validity of the query language $\mathscr{Q}$ is recursively enumerable.) According to Lines 11-16 in Procedure 1, $t_{0}$ will be rendered as true in a finite time.
Let $n>0$. As the inductive hypothesis, we assume that, for every $i\in\{0,\dots,n-1\}$, the task $t_{i}$ has been rendered as true. We need to show that the task $t_{n}$ will also be rendered as true. Only consider Case 3, i.e., the case where $D=D_{j}$ and $\phi_{j}\vDash\phi$ for some $j\in\{0,\dots,n-1\}$. Proofs for other cases are similar, and we omit them here. For Case 3, it is easy to see that $t_{j}=\langle(M_{e},\phi_{j},\phi),(M,D_{j},\phi_{j})$ and $t_{n}$ is the parent task of $t_{j}$ according to Procedure 1. By the inductive hypothesis, the task $t_{j}$ will be rendered as true in a finite time. Through the truth propagation of $F$ (see Lines 14-16 in Procedure 1), we can conclude that $t_{n}$ must also be rendered as true in a finite time.

### v2

@@ Introduction
Knowledge is extensively acknowledged as a cornerstone of intelligence (McCarthy and Hayes 1981; Delgrande et al. 2024), playing a crucial role in intelligent systems. How to effectively represent, acquire, utilize and evolve knowledge is undoubtedly one of the most critical parts of realizing artificial general intelligence (AGI). Knowledge representation serves as the foundation and starting point for all these tasks. In traditional knowledge representation and reasoning (KR), representations of knowledge are often regarded as “explicit, symbolic, declarative representations of information” (Delgrande et al. 2024). However, in this work, we will consider more general forms of knowledge representation.
Over the past seven decades, researchers have devoted substantial efforts to developing various knowledge representation formalisms. An incomplete list includes: monotonic logical systems such as Prolog (van Emden and Kowalski 1976) and description logics (Baader et al. 2017); non-monotonic logics such as circumscription (McCarthy 1980) and default logic (Reiter 1980); graph-based representations such as Bayesian networks (Pearl 1985) and semantic network (Sowa 1991), and parametrized models such as recurrent neural networks (Rumelhart, Hinton, and Williams 1986) and transformers (Vaswani et al. 2017). It is noteworthy that machine learning and knowledge representation are inherently intertwined; all learning algorithms are actually based on some formalisms for representing knowledge.
The quest for the best formalism of knowledge representation has sparked a longstanding dispute. A prime example is the “declarative vs. procedural” controversy, which centers on choosing between declarative statements and procedures as the primary means of representing knowledge. Similarly, the ongoing debate between symbolic AI and connectionist AI revolves around the question of whether knowledge should be represented explicitly or implicitly. In general, knowledge representation formalisms in machine learning, such as convolutional neural networks in deep learning and transformers in large language models, are implicit, while all logical formalisms in traditional KR are explicit.
@@
The main contributions of this work are threefold. Firstly, we propose a general framework to capture all the knowledge representation formalisms in which we are interested, and propose a novel definitions for universal formalisms. Secondly, we find a family of universal formalisms; based on them, we then prove that all possible universal formalisms are recursively isomorphic. Thirdly, we show that, under a natural condition, all subrecursive formalisms that can be translated into each other are recursively isomorphic. These results show us that, if an offline compilation is allowed, almost all the natural representation formalisms with the same expressive power can be regarded as the same, which thus provides us a partial answer to the aforementioned dispute.
@@
Given a subclass $K$ of $\mathscr{D}\times\mathscr{Q}$, let $\textit{cl}(K)$ denote the minimum superclass of $K$ which admits Properties 1-5 of Definition 3. We need to assure the desired transformation $(\cdot)^{\ast}$ satisfying the property: $\mathbb{K}(M^{\ast})=\textit{cl}(\mathbb{K}(M))$.
@@
To address these issues, we introduce a task array $T$ where the $i$-th task is stored in $T[i]$. Although there could be an infinite number of nonterminating tasks in $T$, a standard technique can be applied to sequentially simulate multiple such tasks. This can be visualized by a Cartesian coordinate system where the $x$-axis represents numbers of steps and the $y$-axis denotes task indices. The simulation is actually a procedure to traverse the first quadrant of this coordinate system, which is implemented by Lines 6-9 of Procedure 1.
An atomic task is a tuple $t:=(N,p_{1},\dots,p_{k})$, where $N$ is a Turing machine and $p_{1},\dots,p_{k}$ its parameters. Performing $t$ involves simulating $N$ on the input $[\![p_{1},\dots,p_{k}]\!]$, with $t$ being successful if $N$ accepts. Every task is either an atomic task or a finite sequence of atomic tasks $t:=\langle t_{1},\dots,t_{n}\rangle$. Executing $t$ means sequentially performing the atomic tasks $t_{1},\dots,t_{n}$. The task $t$ is said to be successful or to succeed if all its constituent atomic tasks are successful.
Moreover, let $M_{e}$ denote a Turing machine that accepts $[\![\phi,\psi]\!]$ iff $\phi,\psi$ are a pair of FO-sentences such that $\phi\vDash\psi$. Let $M_{n}$ denote a Turing machine that accepts $[\![\tau]\!]$ iff $\tau$ is a partial injective functions on $\Delta$ with a finite domain.
@@
to $T$, where $\chi$ is the next query after $\psi$ in $\mathscr{Q}$. By repeating this process, we effectively enumerate all queries in $\mathscr{Q}$. For details, see Lines 22-23, 27-28, and 42-47 of Procedure 1.
@@
A flag array $F$ records the completion status of each task as either true or false. The relationships between tasks are tracked using arrays $p$ (parent) and $c$ (cooperation). Specifically, $p[n]=i$ indicates that the task $T[n]$ is generated from the task $T[i]$, making $T[i]$ the parent of $T[n]$. A value of $c[i]=-1$ signifies that the task $T[i]$ operates independently. If $T[i]$ succeeds, its parent task’s flag, $F[p[i]]$, is set to true. Otherwise, if $c[i]=j\neq-1$, the task $T[i]$ must cooperate with the task $T[j]$. Only when both $T[i]$ and $T[j]$ succeed does the parent task’s flag, $F[p[i]]$, get set to true. For details on truth propagation, refer to Lines 14-16 of Procedure 1.
@@ 5 $n\leftarrow 1$; /* There is one task in the current array */
6 for $k\leftarrow 0$ to $\infty$ do
7       for $i\leftarrow 0$ to $\min(n-1,k)$ do
@@ 16                         $F[j]\leftarrow$ true;
17                  if $F[0]$ /* a support found */ then accept;
18
@@ 57                   $n\leftarrow n+1$;
Let $\Theta$ denote the mapping which maps $[\![M^{\ast}]\!]$ to $\mathbb{K}(M^{\ast})$ for every Turing machine $M$.
@@
The recursion theorem can also be generalized to KRFs.
@@
We are now in the position to establish the main theorem.
@@
Every FO-sentence $\phi$ is logical equivalent to $\phi\wedge\phi$. The defined KBs are thus the same, which provides a way to effectively find theories specified to the same KB. Thus, $\Gamma_{\mathrm{FO}}$ (see Example 4) admits the padding property.
@@ Appendix A Appendix: Detailed Proofs
Take $\gamma_{D_{0},\psi}\in\Sigma$ arbitrarily. If there exists no injective assignment $s$ in $\mathcal{B}$ such that $\mathcal{B}\models\theta_{D_{0}}^{v}[s]$, then $\mathcal{B}$ is trivially a model of $\gamma_{D_{0},\psi}$. Otherwise, let $s$ be any of such assignments. Let $\tau$ be a mapping that maps each $c\in\Delta$ to $s(v_{c})$. It is not difficult to verify that $\tau$ is injective and $D$ extends $\tau(D_{0})$. By the construction of $\Sigma$ and $\gamma_{D_{0},\psi}\in\Sigma$, we know $(D_{0},\psi)\in K$. Since $\mathscr{Q}$ is closed under constant renaming, it must be true that $\tau(\psi)\in\mathscr{Q}$. According to Property 5 of Definition 3, we thus have $(\tau(D_{0}),\tau(\psi))\in K$, and consequently, $\tau(\psi)\in\Psi$. This implies that $\mathcal{B}$ is a model of $\tau(\psi)$, or equivalently, $\mathcal{B}\models\psi[s]$. As a consequence, $\mathcal{B}$ is also a model of $\gamma_{D_{0},\psi}$ in this case. By the arbitrariness of $\gamma_{D_{0},\psi}$, we then conclude that $\mathcal{B}$ is a model of $\Sigma$. By definition, it is easy to see that $\mathcal{B}|_{\sigma_{Q}}=\mathcal{A}$. Consequently, we have $\mathcal{A}\models\phi$. This thus proves $\Psi\vDash_{u}\phi$ that we need.
@@
For convenience, we say a task in $T$ is render as true if the flag in $F$ corresponding to the task is set to true.
We first consider the direction of $\subseteq$”. Let $(D,\phi)\in\mathbb{K}(M^{\ast})$. Now our task is to prove that $(D,\phi)\in\textit{cl}(\mathbb{K}(M))$. By definition, $M^{\ast}$ will accept the input $[\![D,\phi]\!]$. Let $N(D,\phi)$ denote the number of iterations of the while-loop (see Lines 14-16 of Procedure 1) before the task in which $(M,D,\phi)$ appears is rendered as true.
@@
We prove this by a routine induction on the length $n$ of $S$. For the case of $n=0$, $S$ consists of exactly one pair $(D,\phi)$, and we have that either $M$ accepts $[\![D,\phi]\!]$ or $\phi$ is a tautology. Thus, $t_{0}=(M,D,\phi)$ if the first case happens, and $t_{0}=(M_{e},\top,\phi)$ otherwise. It is easy to see that in either case performing $t_{0}$ can be done in a finite time. (Note that the validity of the query language $\mathscr{Q}$ is recursively enumerable.) According to Lines 11-16 of Procedure 1, $t_{0}$ will be rendered as true in a finite time.
Let $n>0$. As the inductive hypothesis, we assume that, for every $i\in\{0,\dots,n-1\}$, the task $t_{i}$ has been rendered as true. We need to show that the task $t_{n}$ will also be rendered as true. Only consider Case 3, i.e., the case where $D=D_{j}$ and $\phi_{j}\vDash\phi$ for some $j\in\{0,\dots,n-1\}$. Proofs for other cases are similar, and we omit them here. For Case 3, it is easy to see that $t_{j}=\langle(M_{e},\phi_{j},\phi),(M,D_{j},\phi_{j})$ and $t_{n}$ is the parent task of $t_{j}$ according to Procedure 1. By the inductive hypothesis, the task $t_{j}$ will be rendered as true in a finite time. Through the truth propagation of $F$ (see Lines 14-16 of Procedure 1), we can conclude that $t_{n}$ must also be rendered as true in a finite time.